OpenStudy (anonymous):

in triangle abc; a=14cm b=9cm and c=6cm.... solve the triangle..... how do i do this without angle measurements?

OpenStudy (anonymous):

if i remember correctly it is often more convenient to solve for the greatest angle first the greatest angle will be opposite the largest side, so maybe you would be better of starting with \[\cos(A)=\frac{b^2+c^2-a^2}{2bc}\] we can work through the steps if you like

OpenStudy (anonymous):

ok and i made a typo on my very first line it should be \[a^2=b^2+c^2-2bc\cos(A)\] or \[\cos(A)=\frac{b^2+c^2-a^2}{2bc}\]

OpenStudy (anonymous):

yes please work through it :)

OpenStudy (anonymous):

in your case \(a=14, b=9, c=6\) so you get \[\cos(A)=\frac{9^2+6^2-14^2}{2\times 9\times 6}=-.731\]rounded

OpenStudy (anonymous):

therefore \(A=\cos^{-1}(-.731)=137\) degrees

OpenStudy (anonymous):

i computed in one step here http://www.wolframalpha.com/input/?i=arcos%28%289^2%2B6^2-14^2%29%2F%282*9*6%29%29

OpenStudy (anonymous):

now you know 3 sides and one angle, abandon the law of cosines and use the law of sines for the next angle because it is easier then for the last angle make sure they add up to 180

OpenStudy (anonymous):

btw law of cosines is easy to remember if you remember pythagoras, because it looks almost the same pythagoras for a right triangle \(c^2=a^2+b^2\) if you do not have a right triangle you need so adjust with \[c^2=a^2+b^2-2ab\cos(C)\]

OpenStudy (anonymous):

oh okay thanks! you have been a huge help!!

OpenStudy (anonymous):

yw, you good from here right? with law of sines?

OpenStudy (anonymous):

yeah i get the law of sines